OpenStudy (anonymous):

Convert the decimal to a fraction. Be sure to show all of your work. 0.62

OpenStudy (anonymous):

move decimal over make 62over 100 reduce to 31 over 100

OpenStudy (anonymous):

0.__ =tenth place 0.0__=hundredth place 0.00__= thousandths place etc 0.1 = 1/10 0.01= 1/100 0.001=1/1000